OpenStudy (anonymous):

Inverse of f(x)=-e^-x +3

OpenStudy (anonymous):

switch x with y and solve for y again ...

OpenStudy (anonymous):

ok, what would i do with the exponent? would I square root on both sides?

OpenStudy (anonymous):

\[x=-e ^{-y}+3\]

OpenStudy (anonymous):

take ln

OpenStudy (anonymous):

lnx = y+3 ---- y = ln(x) -3

OpenStudy (anonymous):

oh!!! ok thank you!!!

OpenStudy (anonymous):

I made a mistake sorry ...

OpenStudy (anonymous):

wats up?

OpenStudy (anonymous):

lnx = ln (-e^-y +3)

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

did you get an answer?

OpenStudy (anonymous):

y= -ln(3-x)

OpenStudy (anonymous):

yea i got it.. i just needed a start lol thanks

OpenStudy (anonymous):

yw